Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

Standard Quantity Standard Price
or Rate
Standard Cost
Direct materials 2.50 ounces $ 22.00 per ounce $ 55.00
Direct labor 0.90 hours $ 16.00 per hour 14.40
Variable manufacturing overhead 0.90 hours $ 2.00 per hour 1.80
$ 71.20

During November, the following activity was recorded relative to production of Fludex:

a. Materials purchased, 14,000 ounces at a cost of $289,800.

b. There was no beginning inventory of materials; however, at the end of the month, 4,050 ounces of material remained in ending inventory.

c. The company employs 26 lab technicians to work on the production of Fludex. During November, they worked an average of 150 hours at an average rate of $15.00 per hour.

d. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $5,000.

e. During November, 3,900 good units of Fludex were produced .
Answer #1
Answer
1) Calculation of Materials Price Variance:
Materials price variance = AQ (AP – SP)
14,000 ($20.7 – $22.00)
18,200 F
Working Note:
Actual Price per ounce is 2,89,800/ 14,000 = $20.7
Calculation of Materials Quantity Variance:
Materials quantity variance = SP (AQ – SQ)
22.00 (9,950 – 9,750)
4,400 U
Working Note:
Actual Quantity = 14,000 – 4,050 = 9,950
Standard Quantity = 3,900 * 2.50 = 9,750
Summary
Materials Price Variance: =18,200 F
Materials Quantity Variance:= 4,400 U
Total Material Cost Variance = 18,200 F+4400 U= 13,800 F